OpenStudy (anonymous):

I wanna get done im in a rush I will most certain fan and medal if anyone helps me :)

OpenStudy (anonymous):

4x-8.3=4.5

OpenStudy (agl202):

Ya, Shore...

OpenStudy (anonymous):

@Agl202 okay great my equation is up

OpenStudy (hitaro9):

Okay, so in this type of problem you're trying to figure out what x is. For example, if the problem were x+2 = 4 You could reword it to say "What +2 is 4?" Which would be 2. You can reach this by subtracting 2 from both sides of the equals sign

OpenStudy (hitaro9):

So, doing the same thing to 4x-8.3 =4.5 We can say, add 8.3 to both sides to get 4x = 4.5+8.3 4x = 12.8

OpenStudy (misty1212):

how fast you want to do it?

OpenStudy (hitaro9):

Now we have "What times 4 is 12.8?" We can, using similar logic, divide by 4, to get x = 3.2

OpenStudy (anonymous):

real quick it late im tired

OpenStudy (hitaro9):



OpenStudy (hitaro9):

If your teacher wants you to show the steps, it'd look something like that.

OpenStudy (hitaro9):

Does that make sense? Basic algebra is going to be SUPER important in future math courses, and having a fundamental understanding now is going to help you not struggle in the future.

OpenStudy (raden_zaikaria):